OpenStudy (anonymous):

2. Are 20, 25 and 30 an example of a Pythagorean triplet? Yes or No, Explain your answer and show work 2. Are 20, 25 and 30 an example of a Pythagorean triplet? Yes or No, Explain your answer and show work @Mathematics

OpenStudy (anonymous):

\[20^2+25^2 = 30^2? \] 400 + 625 = 900 no so therefore it is not

OpenStudy (anonymous):

so why not tho?